# Determinantal random subgraphs

### v3

@@ 5.2. Determinantal measures on finite sets
Set $d=\dim E$ and $n=\dim H$. Let $(h_{1},\ldots,h_{n})$ be a basis of $H$. In $^{n}E$, set $\eta=h_{1}\wedge\ldots\wedge h_{n}$. Also, for every subset $I$ of $S$, let us denote by $e_{I}$ the exterior product of the basis elements $\{e_{i}:i\in I\}$. Since $\eta$ is a non-zero element of $^{n}E$, and $\{e_{I}:I\subseteq S,|I|=n\}$ is an orthonormal basis of this space, we have
@@
Proof.
In the language of the exterior algebra, this equality is a consequence of the fact that the Hodge operator on $E$ (see [KL22, Section 5.6]) is an isometry which sends $\eta/\|\eta\|$ to ±$\pm\zeta/\|\zeta\|$ and $e_{I}$ to ±$\pm e_{S\setminus I}$.
@@
[…]
@@ 5.3. A criterion of determinantality
[…]
@@ 5.4. The mean projection theorem
We use the notation $\mathsf{a}$ for the endomorphism of the exterior algebra $E$ of $E$ induced by an operator $\mathsf{a}\in\mathrm{End}(E)$ (see [KL22, Section 5.4]). In the basis $\{e_{I}:I\subseteq S\}$ of $E$, the coefficient $(I,J)$ of the matrix of $\mathsf{a}$ is the $(I,J)$-minor of the matrix of $a$ in the basis $(e_{i})_{i\in S}$.
@@
$\mathbb{E}\big{[}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}\mathsf{P}^{H}_{{\mathsf{X}}^{c}}\big{]}=\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}{\sf\Pi}^{H}\ \text{ and }\ \mathbb{E}\big{[}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}\mathsf{P}^{H^{\perp}}_{{\mathsf{X}}}\big{]}=\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}{\sf\Pi}^{H^{\perp}}.$
@@ 5.5. A conditional probability measure
Let us decompose $\eta$ on the orthonormal basis $\{e_{I}:I\subseteq S,|I|=n\}$ of $^{n}E$ and apply $({\sf\Pi}^{E_{K}})^{\wedge n}$ to both sides of the equality. We find
@@
[…]
@@ 6. The uniform spanning tree
According to (54), in order to understand the distribution of ${\mathsf{X}}$, we need to find the decomposition of the exterior product of the elements of a basis of $\mathop{\mathrm{im}}d$ on the orthogonal basis $\{e^{\star}_{S}:S\subseteq{\sf E}^{+},|S|=|{\sf V}|-1\}$ of $^{|{\sf V}|-1}\Omega^{1}(\mathsf{G})$. This is exactly what is given to us by (45) and we read on this equality, firstly, that ${\mathsf{X}}$ is almost surely a spanning tree of $\mathsf{G}$, and secondly, for all spanning tree $T$, that $\mathbb{P}({\mathsf{X}}=T)\propto\|e^{\star}_{T}\|^{2}=\underline{x}^{T}$, which is the expected result. ∎
@@
The kernel of $d^{*}d$, which is equal to the kernel of $d$, is the line of constant functions on $\mathsf{G}$. Therefore, 00 is a simple eigenvalue of $d^{*}d$, and the determinant of the compression of $d^{*}d$ on the orthogonal of its kernel is equal to the coefficient of the monomial of degree $1$ of its characteristic polynomial. This is the sum of all the minors of size $|{\sf V}|-1$ of its matrix in a basis of $\Omega^{0}(\mathsf{G})$, and we just observed that in the canonical basis, these minors are all equal, namely to $\mathsf{T}(\underline{x})$. ∎
@@
For every integer $k\geqslant 0$, the following equalities of endomorphisms of $^{k}\Omega^{1}(\mathsf{G})$ hold:
@@ 7. Geometric multilinear identities
Let us fix an integer $k\geqslant 0$. For every connected spanning subgraph $K\in\mathcal{C}_{k}(\mathsf{G})$ of $\mathsf{G}$ with $k$ linearly independent cycles, the exterior product $\det\mathscr{Z}^{K}$ of the elements of an integral basis of $Z_{1}(K)$ is defined only up to a sign, but its tensorial square is well defined, and we gave an expression of it in Proposition 3.2. The $k$-th exterior power of the antilinear isomorphism ${\sf J}^{1}_{\underline{x}}$ defined by (33) sends $\det\mathscr{Z}^{K}$ in the inner product space $^{k}\Omega^{1}(\mathsf{G})$, of which the family $(e^{\star}_{S}/\sqrt{\underline{x}^{S}}:S\subseteq{\sf E}^{+},|S|=k)$ is an orthonormal basis.
The following proposition is a geometric variant of Proposition 3.2. The identity is stated in the tensor square over $\mathbb{R}$ of $^{k}\Omega^{1}(\mathsf{G})$, even if the base field is $\mathbb{K}=\mathbb{C}$, in order to deal later with the sesquilinearity of the inner product.
@@
In $^{k}\Omega^{1}(\mathsf{G})\otimes_{\mathbb{R}}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{1}(\mathsf{G})$, the following equality holds:
@@
an equality originally in $^{k}C_{1}(\mathsf{G})\otimes_{\mathbb{Z}}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}C_{1}(\mathsf{G})$, and that still holds in $^{k}\Omega_{1}(\mathsf{G})\otimes_{\mathbb{R}}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega_{1}(\mathsf{G})$ after replacing ${\sf Z}_{T}$ by ${\sf Z}_{T}\otimes\mathrm{id}_{\mathbb{K}}$.
@@
In $^{k}\Omega^{1}(\mathsf{G})\otimes_{\mathbb{R}}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{1}(\mathsf{G})$, the following equality holds:
@@ 8. Determinantal spanning subgraphs
In general, and in contrast with the case of spanning trees, the probability measures that we will define are not uniform. In fact, in view of the complexity results stated below (see Section 15.1), the uniform measure on $\mathcal{C}_{k}(\mathsf{G})$ and $\mathcal{F}_{k}(\mathsf{G})$ cannot be determinantal in general (otherwise, summing over $k$, there would be a determinantal formula for enumerating $\mathcal{C}(\mathsf{G})$ and $\mathcal{F}(\mathsf{G})$, contradicting the #$\#P$-hardness of these counts). Studying the uniform measure would be a more difficult task; on that matter, see [GW04] for conjectures about the uniform measure on connected subgraphs, or spanning forests, without constraint on the Betti numbers.
Figure 2. A random element of $\mathcal{C}_{4}(\mathsf{G})$ (its $2$-core is represented by thickened edges) and a random element of $\mathcal{F}_{4}(\mathsf{G})$ on a hexagonal grid determined by $4$ random $1$-forms (note that one of the trees has zero edges). The sampling algorithm we used here, and for the other figures, is the classical one of [HKPV06, Algorithm 18] for sampling determinantal probability measures.
@@ 8.1. Determinantal connected spanning subgraphs
For every subgraph $K\in\mathcal{C}_{k}(\mathsf{G})$ of $\mathsf{G}$, the $\mathbb{Z}$-module $Z_{1}(K)$ has rank $k$. The exterior product of the elements of an integral basis $\mathscr{Z}^{K}$ of $Z_{1}(K)$ depends on the choice of this basis only up to a sign: we denote it by $\det\mathscr{Z}^{K}$ and see it as an element of $^{k}\Omega_{1}(\mathsf{G})$.
@@
The $1$-forms $\theta_{1},\ldots,\theta_{k}$ still being those used to define $H$, and with the obvious notation, let us consider the basis $\mathscr{H}_{T}=\{d\mathbbm{1}_{v}:v\neq v_{0}\}\cup\{\theta^{\prime}_{1},\ldots,\theta^{\prime}_{k}\}$ of $H$. Let us observe that since $\theta^{\prime}_{i}$ differs from $\theta_{i}$ by a coboundary, the exterior product of the elements of $\mathscr{H}_{T}$ is, up to a sign, equal to that of the elements of =$\mathscr{H}=\{d\mathbbm{1}_{v}:v\neq v_{0}\}\cup\{\theta_{1},\ldots,\theta_{k}\}$. In particular, $\|\det\mathscr{H}_{T}\|^{2}=\|\det\mathscr{H}\|^{2}$ does not depend on $T$.
@@
so that the determinant is equal to ±$\pm\underline{x}^{K\setminus T}(\vartheta,z_{K})$.
@@ 8.2. Determinantal acyclic spanning subgraphs
$\mathscr{H}^{\perp}={\sf J}^{1}_{\underline{x}}(\mathscr{Z})\cup\{\sigma_{1},\ldots,\sigma_{k}\}\ \text{ and }\ \mathscr{H}^{\perp}_{T}={\sf J}^{1}_{\underline{x}}(\mathscr{Z})\cup\{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{k}\}$
@@
so that the last factor is equal to ±$\pm\underline{x}^{T\setminus F}(\det\mathscr{B}^{F},{\sf c})$. Finally, we find
@@ 9.2. The case of two-dimensional simplicial complexes
Figure 5. The $2$-core of a uniform element of $\mathcal{H}_{1}(X)\subset\mathcal{C}_{4}(X_{1})$ where $X=(X_{0},X_{1},X_{2})$ is a discretization of a surface of genus $2$.
Our construction can be generalized to higher dimensional complexes, following [Lyo09], see the brief discussion in Section 14. The above-mentioned duality in the planar case then is the Poincaré duality $X\mapsto X^{\dagger}$ between $k$-cells and $(d-k)$-cells in a $d$-dimensional complex.
@@ 9.3. Choice of convention
Note that we could have given alternative definitions for the polynomials $\mathsf{A}_{\mathsf{G}}(\vartheta,\underline{x})$, replacing the terms $\underline{x}^{F}$ by $\underline{x}^{F^{c}}$, like in the definition of Symanzik polynomials (Section 10.4). This would have had the advantage of simplifying certain formulas, notably those that make use of the planar duality, such as (73). However, we have chosen to endow only cochains (that is, 00-forms and $1$-forms) with inner products (see Section 4.4) and we also prefer to define determinantal processes on the set of edges with respect to subspaces of $\Omega^{1}(\mathsf{G})$ (not of $\Omega_{1}(\mathsf{G})$), so as to compare more easily with the classical cases of the uniform spanning tree. Defining the polynomials so as they would be generating functions for these determinantal probability measures (and not their dual determinantal probability measures) was a further argument in favor of this choice. This choice will also be apparent in the way we treat with the matroid generalisation in Section 13.
@@ 10.2. Symanzik and Kirchhoff polynomials
[…]
@@ 11.1. Intersection of hyperplanes and random spanning forests
Since $d$ is $1$-to-$1$ on $\mathop{\mathrm{im}}d^{*}$ and since $H_{{\sf q}}=\{f\in\mathop{\mathrm{im}}d^{*}:\forall i\in\{1,\ldots,k\},\langle f,q_{i}\rangle=0\}$ is an intersection of $k$ hyperplanes, the codimension of $H_{{\sf q}}$ in $\mathop{\mathrm{im}}d$ is $\dim{\mathrm{Vect}}(q_{1},\ldots,q_{k})$. ∎

### v4

@@ 5.2. Determinantal measures on finite sets
[…]
@@
$|\langle\eta,e_{I}\rangle|^{2}/ \|\eta\|^{2}=|\langle\zeta,e_{S\setminus I}\rangle|^{2}/\|\zeta\|^{2}.$
Proof.
@@
$|\langle\eta,e_{I}\rangle|^{2}/ \|\eta\|^{2}=|\det(A^{I})|^{2}/\det(A^{\dagger}A)$ $=|\det(R^{I}_{\{1,\ldots,n\}})|^{2}$
$\hskip-56.9055pt=|\det(R^{S\setminus I}_{[d]\setminus[n]})|^{2}=|\det(B^{I})|^{2}/\det(B^{\dagger}B)=|\langle\zeta,e_{S\setminus I}\rangle|^{2}/ \|\zeta\|^{2}$
@@ 5.3. A criterion of determinantality
For endomorphisms defined on $E$ or with values in $E$, we will denote by $\dagger$ the adjunction with respect to the original inner product $\langle\cdot,\cdot\rangle$ and by * the adjunction with respect to the twisted inner product $\langle\cdot,\cdot\rangle_{\underline{x}}$. It should therefore be borne in mind that an expression such as $a^{*}$ depends on $\underline{x}$.
@@ 5.4. The mean projection theorem
[…]
@@
[…]
@@ 5.5. A conditional probability measure
[…]
@@
$\mathbb{P}({\mathsf{X}}=I| {\mathsf{X}}\subseteq K)=\frac{|\langle\eta,e_{I}\rangle|^{2}}{\|\eta\|^{2}}\frac{\|\eta\|^{2}}{\|\phi\|^{2}}=\frac{|\langle\eta,e_{I}\rangle|^{2}}{\|\phi\|^{2}}=\frac{|\langle\phi,e_{I}\rangle|^{2}}{\|\phi\|^{2}},$
@@ 6. The uniform spanning tree
According to (54), in order to understand the distribution of $\mathsf{X}$}, we need to find the decomposition of the exterior product of the elements of a basis of $\mathop{\mathrm{im}}d$ on the orthogonal basis $\{e^{\star}_{S}:S\subseteq{\sf E}^{+},|S|=|{\sf V}|-1\}$ of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{|{\sf V}|-1}\Omega^{1}(\mathsf{G})$. This is exactly what is given to us by (45) and we read on this equality, firstly, that $\mathsf{X}$} is almost surely a spanning tree of $\mathsf{G}$, and secondly, for all spanning tree $T$, that $\mathbb{P}({\mathsf{X}}=T)\propto\|e^{\star}_{T}\|^{2}=\underline{x}^{T}$, which is the expected result. ∎
@@
The kernel of $d^{*}d$, which is equal to the kernel of $d$, is the line of constant functions on $\mathsf{G}$. Therefore, 0 is a simple eigenvalue of $d^{*}d$, and the determinant of the compression of $d^{*}d$ on the orthogonal of its kernel is equal to the coefficient of the monomial of degree $1$ of its characteristic polynomial. This is the sum of all the minors of size $|{\sf V}|-1$ of its matrix in a basis of $\Omega^{0}(\mathsf{G})$, and we just observed that in the canonical basis, these minors are all equal, namely to $\mathsf{T}(\underline{x})$. ∎
@@
For every integer $k\geqslant 0$, the following equalities of endomorphisms of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{1}(\mathsf{G})$ hold:
@@ 7. Geometric multilinear identities
Let us fix an integer $k\geqslant 0$. For every connected spanning subgraph $K\in\mathcal{C}_{k}(\mathsf{G})$ of $\mathsf{G}$ with $k$ linearly independent cycles, the exterior product $\det\mathscr{Z}^{K}$ of the elements of an integral basis of $Z_{1}(K)$ is defined only up to a sign, but its tensorial square is well defined, and we gave an expression of it in Proposition 3.2. The $k$-th exterior power of the antilinear isomorphism ${\sf J}^{1}_{\underline{x}}$ defined by (33) sends $\det\mathscr{Z}^{K}$ in the inner product space $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{1}(\mathsf{G})$, of which the family $(e^{\star}_{S}/\sqrt{\underline{x}^{S}}:S\subseteq{\sf E}^{+},|S|=k)$ is an orthonormal basis.
The following proposition is a geometric variant of Proposition 3.2. The identity is stated in the tensor square over $\mathbb{R}$ of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{1}(\mathsf{G})$, even if the base field is $\mathbb{K}=\mathbb{C}$, in order to deal later with the sesquilinearity of the inner product.
@@
[…]
@@
[…]
@@
[…]
@@ 8. Determinantal spanning subgraphs
In general, and in contrast with the case of spanning trees, the probability measures that we will define are not uniform. In fact, in view of the complexity results stated below (see Section 15.1), the uniform measure on $\mathcal{C}_{k}(\mathsf{G})$ and $\mathcal{F}_{k}(\mathsf{G})$ cannot be determinantal in general (otherwise, summing over $k$, there would be a determinantal formula for enumerating $\mathcal{C}(\mathsf{G})$ and $\mathcal{F}(\mathsf{G})$, contradicting the #P\#P-hardness of these counts). Studying the uniform measure would be a more difficult task; on that matter, see [GW04] for conjectures about the uniform measure on connected subgraphs, or spanning forests, without constraint on the Betti numbers.
Figure 2. A random element of $\mathcal{C}_{4}(\mathsf{G})$ (its 22-core is represented by thickened edges) and a random element of $\mathcal{F}_{4}(\mathsf{G})$ on a hexagonal grid determined by 44 random 11-forms (note that one of the trees has zero edges). The sampling algorithm we used here, and for the other figures, is the classical one of [HKPV06, Algorithm 18] for sampling determinantal probability measures.
@@ 8.1. Determinantal connected spanning subgraphs
For every subgraph $K\in\mathcal{C}_{k}(\mathsf{G})$ of $\mathsf{G}$, the $\mathbb{Z}$-module $Z_{1}(K)$ has rank $k$. The exterior product of the elements of an integral basis $\mathscr{Z}^{K}$ of $Z_{1}(K)$ depends on the choice of this basis only up to a sign: we denote it by $\det\mathscr{Z}^{K}$ and see it as an element of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega_{1}(\mathsf{G})$.
@@
The $1$-forms $\theta_{1},\ldots,\theta_{k}$ still being those used to define $H$, and with the obvious notation, let us consider the basis $\mathscr{H}_{T}=\{d\mathbbm{1}_{v}:v\neq v_{0}\}\cup\{\theta^{\prime}_{1},\ldots,\theta^{\prime}_{k}\}$ of $H$. Let us observe that since $\theta^{\prime}_{i}$ differs from $\theta_{i}$ by a coboundary, the exterior product of the elements of $\mathscr{H}_{T}$ is, up to a sign, equal to that of the elements of $\mathscr{H}=\{d\mathbbm{1}_{v}:v\neq v_{0}\}\cup\{\theta_{1},\ldots,\theta_{k}\}$. In particular, $\|\det\mathscr{H}_{T}\|^{2}=\|\det\mathscr{H}\|^{2}$ does not depend on $T$.
@@
so that the determinant is equal to $\pm\underline{x}^{K\setminus T}(\vartheta,z_{K})$.
@@ 8.2. Determinantal acyclic spanning subgraphs
[…]
@@
so that the last factor is equal to $\pm\underline{x}^{T\setminus F}(\det\mathscr{B}^{F},{\sf c})$. Finally, we find
@@ 9.2. The case of two-dimensional simplicial complexes
Figure 5. The 22-core of a uniform element of $\mathcal{H}_{1}(X)\subset\mathcal{C}_{4}(X_{1})$ where $X=(X_{0},X_{1},X_{2})$ is a discretization of a surface of genus $2$.
Our construction can be generalized to higher dimensional complexes, following [Lyo09], see the brief discussion in Section 14. The above-mentioned duality in the planar case then is the Poincaré duality $X\mapsto X^{\dagger}$ between $k$-cells and (dk)(d-k)-cells in a $d$-dimensional complex.
@@ 9.3. Choice of convention
Note that we could have given alternative definitions for the polynomials $\mathsf{A}_{\mathsf{G}}(\vartheta,\underline{x})$, replacing the terms $\underline{x}^{F}$ by $\underline{x}^{F^{c}}$, like in the definition of Symanzik polynomials (Section 10.4). This would have had the advantage of simplifying certain formulas, notably those that make use of the planar duality, such as (73). However, we have chosen to endow only cochains (that is, 0-forms and $1$-forms) with inner products (see Section 4.4) and we also prefer to define determinantal processes on the set of edges with respect to subspaces of $\Omega^{1}(\mathsf{G})$ (not of $\Omega_{1}(\mathsf{G})$), so as to compare more easily with the classical cases of the uniform spanning tree. Defining the polynomials so as they would be generating functions for these determinantal probability measures (and not their dual determinantal probability measures) was a further argument in favor of this choice. This choice will also be apparent in the way we treat with the matroid generalisation in Section 13.
@@ 10.2. Symanzik and Kirchhoff polynomials
and $\langle\cdot,\cdot\rangle$ is a Minkowski bilinear form on $\mathbb{R}^{D}$. Here we used the notation
@@ 11.1. Intersection of hyperplanes and random spanning forests
Since $d$ is 11-to-11 on $\mathop{\mathrm{im}}d^{*}$ and since $H_{{\sf q}}=\{f\in\mathop{\mathrm{im}}d^{*}:\forall i\in\{1,\ldots,k\},\langle f,q_{i}\rangle=0\}$ is an intersection of $k$ hyperplanes, the codimension of $H_{{\sf q}}$ in $\mathop{\mathrm{im}}d$ is $\dim{\mathrm{Vect}}(q_{1},\ldots,q_{k})$. ∎